### v3

# VocSegMRI: Multimodal Learning for Precise Vocal Tract Segmentation in Real-time MRI

Abstract
Accurately segmenting articulatory structures in real-time magnetic resonance imaging (rtMRI) remains challenging, as most existing methods rely almost entirely on visual cues. Yet synchronized acoustic and phonological signals provide complementary context that can enrich visual information and improve precision. In this paper, we introduce VocSegMRI, a multimodal framework that integrates video, audio, and phonological inputs through cross-attention fusion for dynamic feature alignment. To further enhance cross-modal representation, we incorporate a contrastive learning objective that improves segmentation performance even when the audio modality is unavailable at inference. Evaluated on a sub-set of USC-75 rtMRI dataset, our approach achieves state-of-the-art performance, with a Dice score of 0.95 and a 95th percentile Hausdorff Distance (HD95) of 4.20 mm, outperforming both unimodal and multimodal baselines. Ablation studies confirm the contributions of cross-attention and contrastive learning to segmentation precision and robustness. These results highlight the value of integrative multimodal modeling for accurate vocal tract analysis.
Index Terms—  Segmentation, Multimodal Learning, Real-time MRI, Vocal Tract
1 Introduction
Fig. 1: Overview of the proposed VocSegMRI model with cross-attention fusion and contrastive supervision.
Real-time magnetic resonance imaging (rtMRI) is increasingly used in speech research, offering a non-invasive, high-temporal-resolution view of the entire vocal tract during continuous speech [1]. Accurate segmentation of vocal tract structures is essential not only for phonetic and linguistic studies but also for clinical applications such as pre-surgical planning in glossectomy and monitoring articulatory decline in Parkinson’s disease [2, 3, 4]. Beyond imaging, acoustic signals and phonological class features capture structured articulatory properties like place and manner of articulation, offering symbolic context that can be integrated with rtMRI for more precise analysis [5, 6].
Early efforts to extract articulatory contours largely relied on manual or semi-automated boundary tracing. A typical pipeline involves hand-annotating air–tissue boundaries in a reference frame, followed by nonlinear optimization propagating these contours across subsequent frames [7]. Other approaches assigned labels by examining pixel intensities along predefined gridlines overlaid on MR images, or by employing active shape models to capture the expected anatomical contours [8]. This process is not only time-consuming and labor-intensive but also susceptible to error, requiring extensive human supervision. More recently, segmentation approaches leveraging deep learning have been proposed [9, 10, 11, 12]. In particular,  [9] employed Fullly Convolutional Networks (FCNs) to label air–tissue boundaries, with [9] additionally assigning boundary pixels to specific articulators. Other works, such as  [10], utilized FCNs resembling the original FCN [13] and the variant in [14], respectively, to annotate air–tissue interfaces. Matthieu Ruthven et al. developed a U-Net-based segmentation model identifying six vocal tract structures, achieving a Dice coefficient of 0.85 [11]. More recently, a multimodal system combining rtMRI with synchronized speech audio was introduced, which uses a fully convolutional architecture alongside Transformer-based fusion, setting a new benchmark in speaker-independent vocal tract segmentation [12].
In this work, we propose VocSegMRI, a multimodal framework for articulator segmentation in rtMRI that integrates visual, acoustic, and phonological inputs. Our contributions are threefold: (i) a framework that incorporates a cross-attention fusion mechanism, enabling the visual encoder to focus on complementary information from the audio and phoneme streams; (ii) a dual-level contrastive learning objective, applied globally and locally, to improve cross-modal alignment and consistency; and (iii) a systematic evaluation on the USC-75 dataset, where our model achieves a Dice coefficient of 0.95 and an HD95 of 4.20 mm, outperforming strong unimodal and multimodal baselines. These results demonstrate the effectiveness of integrative multimodal modeling for more precise vocal tract segmentation.
2 Materials and Methods
2.1 Data
We utilized data from five participants (one male and four females) from the USC-75 dataset for training and in-domain evaluation. Each participant was instructed to read the Rainbow and North Wind and the Sun passages [15, 16]. The real-time MRI data were acquired at 83.28 fps with 2.4 mm in-plane resolution ($84\times 84$ pixels) and 6 mm slice thickness [17]. Imaging used a 1.5 T GE Signa Excite scanner; synchronized audio was recorded at 20 kHz via a fiber-optic microphone, hardware-locked to the scanner clock and denoised. The binary segmentation masks were manually annotated and further reviewed by a speech therapist expert, with recognized expertise in the field. Phonemes were obtained from the outputs of a phonological classifier, aligned with the corresponding phoneme labels from the audio stream, and subsequently refined through manual correction [5, 6].
To expand the training set, data augmentation was applied to each image using a combination of geometric and intensity-based transformations. Augmented images were subsequently resized to a resolution of $224\times 224$ pixels. Each dataset was individually normalized based on its minimum and maximum intensity values to ensure pixel intensities were within the range [0, 1]. As a result, a total of 14,406 usable images were obtained for training and evaluation.
2.2 VocSegMRI
The overall framework of our proposed model is illustrated in Fig. 1. It integrates visual, acoustic, and phonological information within a cross-attention Transformer architecture, complemented by contrastive learning. In the encoder stage, a pretrained Vision Transformer (ViT) based on the google/vit-base-patch16-224-in21k checkpoint, extracts spatial representations from rtMRI frames [18, 19], while synchronized audio is encoded with a pretrained WavLM model from the microsoft/wavlm-base-plus checkpoint [20], and phonological features are mapped using a lightweight MLP. The latter two are combined and projected to form multimodal memory tokens, which interact with image queries through cross-attention layers in the Transformer decoder [21], enabling modality-aware integration for refined segmentation. To further strengthen alignment across modalities, a contrastive module projects image, audio, and phonological tokens into a shared latent space. This supervision encourages the model to learn more discriminative visual features by leveraging audio–phonological guidance. As a result, the model supports inference using video alone, without requiring additional modalities. The final training objective integrates cross-entropy, Dice, and contrastive losses. The source code will be released upon acceptance.
3 Experiments and results
Among the five available participants, we adopted a leave-one-speaker-out strategy. To ensure fair comparisons, all experiments were conducted under identical preprocessing pipelines and without any postprocessing. During training, the audio encoder was frozen to retain its pretrained acoustic representations, while the ViT encoder, which was initially frozen and progressively unfrozen to allow gradual fine-tuning for adaptation to the rtMRI domain. All models were optimized using AdamW with a learning rate of $1\mathrm{e}-4$ and a batch size of 16, trained on an NVIDIA RTX A100 GPU. Training until convergence on the validation set, with early stopping applied (patience=15) to prevent overfitting.
We designed three experimental groups, with results presented in the upper, middle, and lower sections of Tab. Evaluation metrics include Intersection over Union (IoU), Dice coefficient, Average Symmetric Surface Distance (ASSD) and HD95. All metrics are reported as mean $\pm$ standard deviation (std). In the Image-Only Baseline, we compared the performance of several state-of-the-art (SOTA) segmentation models using only frame-wise rtMRI images as input [22, 23, 24, 25, 26, 27]. Among them, ViT-base and nnU-Net achieved the highest IoU of 0.86, followed by ResNet (0.84) and SAM-Med2D (0.83). In the Concat Fusion setting, we implemented multimodal fusion by concatenating the outputs of the respective encoders for each modality. We evaluated segmentation performance using different combinations of input modalities. Adding either audio or phonological class features led to improvements in IoU over the ViT (0.86), with varying magnitudes. The best performance was achieved using all three modalities (IoU==0.89). Finally, in the Ablation Study configuration, we evaluated the contribution of each component individually. The Cross configuration incorporates only cross-attention, while Contrast introduces only contrastive learning. Our proposed model, VocSegMRI, combines both mechanisms, achieving the best overall performance. yielding the highest Dice score (0.95) and the lowest ASSD (1.52) and HD95 (4.26).
a
b
Fig. 2: Per-class segmentation performance of the proposed VocSegMRI model on the test set. (a) Dice coefficient. (b) 95th percentile Hausdorff Distance (HD95).
Fig. 2 presents the per-class segmentation performance of the proposed VocSegMRI model on the test set. As shown in Fig. 2(a), the model achieved the highest Dice scores for the Tongue and Velum classes, with median values exceeding 0.95 and 0.93, respectively, and narrow interquartile ranges, indicating stable segmentation performance. In contrast, the Upper Lip and Lower Lip exhibited greater variability and generally lower Dice scores, with medians around 0.85 and 0.83, respectively. A similar trend is observed in Fig. 2(b), the median HD95 for the Lower Lip class exceeds 4 mm, with a wide distribution and many outliers.
We selected four representative models for comparison. Fig presents qualitative segmentation results for the Tongue and Lower Lip, highlighting false positive (FP) and false negative (FN) regions, as well as reporting precision/recall values, while nnU-Net performed the worst among all models. Specifically, for the Tongue, nnU-Net achieved a precision of 0.91 and recall of 0.80, while for the Lower Lip, precision was 0.42 and recall 0.97. Incorporating all three modalities as input led to improved model performance. However, VocSegMRI consistently achieved the best trade-off across both structures, with an overall precision of 0.85 and recall of 0.98.
4 Discussion and Conclusions
The experimental results demonstrate that our proposed multimodal segmentation framework, VocSegMRI, effectively enhances segmentation performance and exhibits superior segmentation precision compared to established baselines. When using only frame-wise rtMRI images as input, the choice of encoder influenced performance, with ViT achieving the highest IoU among single-modality models. This validates the effectiveness of Transformer-based visual feature extraction for anatomical structures.
In the Concat Fusion setting, we investigated the impact of incorporating additional modalities. Concatenating audio or phonological features with visual information led to improvements over the video-only baseline (IoU 0.86), with the best results obtained when all modalities were used simultaneously (IoU 0.89). Notably, adding phonological features alone provided a smaller performance gain compared to incorporating audio signals. This limited improvement may be attributed to the difficulty of learning meaningful mappings between phonological labels and visual articulator regions, which reduces their ability to guide the model’s attention to specific areas. Overall, these results confirm that multimodal inputs provide complementary information, enhancing segmentation quality by supplying both dynamic articulatory cues and structured phonological information.
The Ablation Study explored advanced fusion mechanisms and contrastive supervision. Replacing simple concatenation with cross-attention allowed the image queries to selectively integrate complementary information from audio and phonological streams, while the addition of a contrastive learning objective further aligned representations across modalities. As a result, VocSegMRI achieved the best overall performance, with a Dice scores of 0.95, ASSD of 1.52 mm, and HD95 of 4.26 mm, surpassing both unimodal and simple concatenation baselines. Notably, the model also exhibited low standard deviations across all metrics, indicating that these gains were consistent across different speakers and highlighting the robustness of the framework.
Class-level analysis (Fig. 2) indicates that segmentation is more accurate for larger structures such as the Tongue and Velum, whereas smaller articulators (Upper/Lower Lips) remain challenging due to their very low pixel representation in the images and weaker visual cues. For instance, in the example shown in Fig, the Lower Lip achieves only 0.14 precision with nnU-Net and 0.42 precision with the ViT baseline, despite high recall values (0.38 and 0.97, respectively), reflecting numerous FP. In contrast, the Tongue exhibits higher precision across models (0.75–0.91) with high recall. Qualitative comparisons further illustrate that unimodal models suffer from pronounced FP and FN, particularly along anatomical boundaries. While simple concatenation partially mitigates these errors, cross-attention combined with contrastive learning enables more precise and balanced segmentation, improving Lower Lip precision to 0.68 and overall precision-recall trade-off to 0.85/0.98, thereby substantially reducing both false positives and false negatives across all structures.
In summary, we proposed a novel multimodal segmentation framework for rtMRI, integrating visual, acoustic, and phonological information through cross-attention fusion and contrastive supervision. Experiments on the USC-75 dataset demonstrate SOTA performance, confirming the effectiveness of modality-aware alignment. The contrastive component ensures reliable segmentation even when speech information is degraded or absent, as in glossectomy patients. Despite these advances, accurate segmentation of small, low-contrast structures remains challenging. Future work will explore adaptive attention mechanisms, temporal modeling, targeted data augmentation, and domain generalization strategies to improve performance on underrepresented articulators and unseen speakers, paving the way for robust, speaker-independent vocal tract analysis.